OpenStudy (anonymous):

the frequency of oscillation of a mass m suspended bu a spring is v1. if the lenght of spring is cut to one third then the same mass oscillates with frequency v2 then relation between v1 and v2is?

OpenStudy (cybershadow):

see frequency of string is f = n(v/2l)

OpenStudy (cybershadow):

now u can take l1 as = l and l2 = l/2

OpenStudy (cybershadow):

therefore v1 = n (v/2l) and v2 = n(v/2(l/2))

OpenStudy (cybershadow):

so v1/v2 = l/2l = 1/2

OpenStudy (cybershadow):

and v1 = 2v2 , Hope its clear :)

OpenStudy (vincent-lyon.fr):

@CyberShadow I think this question is not about compression or transverse waves travelling on a massive spring. It is just a massless spring with a mass attached to it. When you cut the length of spring, you change the overall spring constant. This is why the mass-spring system will oscillate at a different frequency.

OpenStudy (cybershadow):

ah Yes! Nevermind

OpenStudy (cybershadow):

@Vincent-Lyon.Fr

OpenStudy (cybershadow):

i missread Spring as STRING

OpenStudy (vincent-lyon.fr):

lol! It sometimes happens to me too!

OpenStudy (anonymous):

If spring force constant remains the same, F = - k x and if we are ignoring the mass of the spring, as usual, then the velocity will be unchanged.